OpenStudy (anonymous):

3y^2+ny+25. Can you please solve for n?

OpenStudy (solomonzelman):

is your question, 3y^2+ny+25=0 . Can you please solve for n?

OpenStudy (anonymous):

Yes, it is the same one.

OpenStudy (solomonzelman):

\(\Large\color{blue}{ \bf 3y^2+ny+25=0}\) I can't really solve this, because it's 1 equation with 2 unknown variables.

OpenStudy (solomonzelman):

\(\Large\color{blue}{ \bf 3y^2+ny+25=0}\) Well IF it has to be factor-able, AND IF it has t be factor-able into natural numbers \(\Large\color{blue}{ \bf (3y+ 5)(y+5)=0}\) \(\Large\color{blue}{ \bf n≥20}\)